OpenStudy (anonymous):

Identify the Exponent Law - (a^2)^3 = a^2·3 = a^6 I think it's power of a power but all of the powers throw me off.

OpenStudy (zehanz):

Power of power law: \[(a^b)^c = a^{bc}\]Why does this hold?\[(a^b)^c=a^b \cdot a^b \cdot a^b \cdot ... \cdot a^b=a^{b+b+b+...+b}=a^{b \cdot c}\]So it makes use of another power law...

OpenStudy (anonymous):

Could it be product of powers?

OpenStudy (zehanz):

To clarify a bit more: in my explanation above, there are c factors a^b. I then used the rule \[a^b \cdot a^c=a^{b+c}\]only with c times the same power, so you end up with c times an exponent b.

OpenStudy (zehanz):

So yes, it is power of powers, but that rule uses product of powers itself ;)

OpenStudy (zehanz):

Put differently: the power of powers rule is a consequence of the product of powers rule, which is a consequence of the definition of powers...

OpenStudy (anonymous):

Still a bit confusing, but I got it! Thank you :)

OpenStudy (zehanz):

Yes, I know, it sounds confusing, but if you look into it, you'll see that it is quite simple, actually! (It's all about counting the factors of a power).